OpenStudy (anonymous):

x goes into 4x how many times?

OpenStudy (amistre64):

divide

OpenStudy (anonymous):

its actually from cross multiplying x is a numerator and 4x is a denominator

OpenStudy (amistre64):

if you have 4 xs, how many xs do you have?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

how many 3s are in 12? 3+3+3+3 = 12 so 4 3s how many xs are in 4x? x+x+x+x = 4x

OpenStudy (anonymous):

x is over 2 and 4x is under 4 i need to cross multiply and cancel

OpenStudy (amistre64):

x/2 = 4x/4 when x/2 = x which is never so 0

OpenStudy (amistre64):

4/4x

OpenStudy (amistre64):

soo, x/2 = 4/4x is your question

OpenStudy (amistre64):

cross multiply as you suggested 4x^2 = 8

OpenStudy (anonymous):

yess okay so it becomes squared?

OpenStudy (amistre64):

4xx = 4x^2 yes

OpenStudy (anonymous):

You should definitely try mathway.... I'm not sponsored I just love the site. try 4 divided by x.

OpenStudy (anonymous):

the answer is 4/x

OpenStudy (anonymous):

Im so sorry, the final answer is 1/4